Find the equation to the straight line which passes through the point 5,6 and has intercept on the axis equal in magnitude and both sides are positive

Grade:10

2 Answers

Aditya Gupta
4 years ago
hello samaira, the equation would be
y – 6= m(x – 5), where m is the slope.
note that its intercepts are (5 – 6/m, 0) and (0, 6 – 5m)
given 5 – 6/m= 6 – 5m
or m= – 1, 6/5.
when m= 6/5, the intercepts are 0. but they should be positive. so m= – 1
hence eqn is 
x + y= 11
kindly approve :)
Vikas TU
4 years ago
 
Dear student 
 
if intercept are equal and opposite in sign then its slope whould be – 1
 
hence equation is x/ a – y/a = 1
 
5/a – 6/a = 1
 
hence a = – 1
 
hence equation is-
 
- x + y = 1